Connect English with Mathematics
1. Translate each sentence into an equation. Tell how you are assigning the two variables in each.
c) Twice a number, decreased by nine, is half the same number, increased by six.
Thanks.

Answer by checkley77(12844)

2X-9=(X/2)+6
2X-X/2=6+9
(2X*2-X)/2=15
(4X-X)/2=15
3X/2=15
3X=15*2
3X=30
X=30/3
X=10 ANS.
PROOF:
2*10-9=10/2+6
20-9=5+6
11=11